OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (1, 8); m = - 1/6

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y-8=-1/6(x-1) and that's in point-slope form